OpenStudy (anonymous):

A hyperbola has vertices (0,-6) and (0,6) and asymptotes y=3/4x and y=-3/4x. The center is (0,0). In what direction is this hyperbola oriented?

OpenStudy (anonymous):

Center of the hyperbola is the midpoint of the vertices. Midpoint of (0,-6) and (0,6) is (0,0) The transverse axis is the x-axis. Equation is x^2 / a^2 - y^2 / b^2 = 0 Asymptote = y^2 = (b^2/a^2) x^2 +/- (b/a) b/a = 3/4 b=3 a=4 x^2 / 16 - y^2 / 9 = 1

OpenStudy (anonymous):

thank you very much! :)